Please help me!! Here is the problem: 3k=-9/5. Thanks
Answer by ilana(307)

We need to isolate k, so we divide both sides by 3. So we get k=(-9/5)/3, or k=(-9/5)*(1/3), so k=(-9*1)/(5*3)=-9/15=-3/5.